# Multi-View Adaptive Contrastive Learning for Information Retrieval Based Fault Localization

### v1

@@ 1 Introduction
Inspired by the recent success of contrastive learning (CL) technique, in this work, we propose Multi-View Adaptive Contrastive Learning (MACL-IRFL) fault localization method to address the above limitations. First, the bug reports, source code files, and the relationships between them, are modeled as structured graphs from three views (i.e., report-code interaction view, report-report similarity view, and code-code co-citation view), and perform message aggregation separately. Then, the contrastive loss forces the representations of a report across report-report view and report-code view (or a code file across code-code view and report-code view) to be closer to each other, while those of different reports (or code files) apart. Ideally, with the help of contrastive learning, the report representations (or code representations) will only encode the information shared between report-report and report-code views (or code-code and report-code views), thereby alleviating the noise of auxiliary information.
@@ 3.1 Motivating examples
Scenario 1. Fig. 2 displays bug reports #$\#56029$ and #$\#33356$, and their relevant buggy files in ground truth.
@@ 5.1.1 Dataset
Birt 06/2005-12/2013 4178 (4009) 8799 2468 28.05%
Eclipse* 10/2001-01/2014 6495 (6261) 13327 2068 15.52%
JDT 10/2001-01/2014 6274 (6108) 14167 1892 13.35%
SWT 02/2002-01/2014 4151 (4086) 3587 768 21.41%
Tomcat 07/2002-01/2014 1056 (1020) 2755 625 22.69%
@@
Birt 7880 35525 12254 1.97 8.86 4.97
Eclipse* 8455 104756 15552 1.35 16.73 7.52
JDT 9608 82450 15238 1.57 13.50 8.05
SWT 6697 150964 4984 1.64 36.95 6.49
Tomcat 1647 4911 1940 1.61 4.81 3.10
@@ 5.3 Experimental Results for Research Questions
Among all compared baselines, our approach achieves the best performance in most cases across all those projects. The main reasons are three folds: 1) MACL-IRFL constructs a multi-view framework that explicitly encodes the intricate interaction features between bug reports and source code files into their respective representation learning processes, which achieves a deep integration of information and gives higher ranking scores. 2) MACL-IRFL does not pay much attention on the initial natural language text in bug reports and programming codes in source files since we assign random initial features to the bug reports and the source code files, which could mitigate the problem of low-quality bug reports and noisy programming codes. 3) MACL-IRFL incorporates and encodes the similarity relationships and the co-citation relationships into the representation learnings of bug reports and source code files with weights, which helps to alleviate the insufficient bug-fix records issue and improve the accuracy of localization.
@@
To investigate whether the high-layer embedding propagation can help improve the performance of MACL-IRFL, we vary the number of layers $L$ of MACL-IRFL from 1 to 5 and compare their performance under different layers. Table 7 shows the overall performance of the five projects, and MACL-IRFL-$L$ denotes our model with $L$ layers embedding propagation. Fig. 9 illustrates the detailed performance of tne five projects at different layer settings. We have the following observations.

### v2

@@ 1 Introduction
Inspired by the recent success of contrastive learning (CL) technique, in this work, we propose Multi-View Adaptive Contrastive Learning (MACL-IRFL) fault localization method to address the above limitations. To tackle the Absence of bug-fixing history information during the prediction stage, we model bug reports, source code files, and the relationships between them as structured graphs from three views (i.e., report-code interaction view, report-report similarity view, and code-code co-citation view). These views provide auxiliary information that supplements the missing historical connections in the validation and test stages. By performing message aggregation within each view, our approach effectively captures meaningful interactions across different types of data. For the Auxiliary information overload challenge, we leverage the power of contrastive learning. Specifically, the contrastive loss aligns the representations of a report across report-report view and report-code view (or a code file across code-code view and report-code view) to be closer to each other, while those of different reports (or code files) apart. This alignment enables the representations of bug reports and source code files to focus on the shared and relevant information between views while filtering out noise from irrelevant auxiliary data.
@@ 3.1 Motivating examples
Scenario 1.
Fig. 2 displays bug reports #$\#56029$ and #$\#33356$, and their relevant buggy files in ground truth.
@@ 5.1.1 Dataset
Birt 06/2005-12/2013 4,178 (4,009) 8,799 2,468 28.05%
Eclipse* 10/2001-01/2014 6,495 (6,261) 13,327 2,068 15.52%
JDT 10/2001-01/2014 6,274 (6,108) 14,167 1,892 13.35%
SWT 02/2002-01/2014 4,151 (4,086) 3,587 768 21.41%
Tomcat 07/2002-01/2014 1,056 (1,020) 2,755 625 22.69%
@@
Birt 7,880 35,525 12,254 1.97 8.86 4.97
Eclipse* 8,455 104,756 15,552 1.35 16.73 7.52
JDT 9,608 82,450 15,238 1.57 13.50 8.05
SWT 6,697 150,964 4,984 1.64 36.95 6.49
Tomcat 1,647 4,911 1,940 1.61 4.81 3.10
@@ 5.3 Experimental Results for Research Questions
Among all compared baselines, our approach achieves the best performance in most cases across all those projects. The main reasons are three folds: 1) MACL-IRFL constructs a multi-view framework that explicitly encodes the intricate interaction features between bug reports and source code files into their respective representation learning processes, which achieves a deep integration of information and gives higher ranking scores. 2) MACL-IRFL does not pay much attention on the initial natural language text in bug reports and programming codes in source files since we assign random initial features to the bug reports and the source code files, which could mitigate the problem of low-quality bug reports and noisy programming codes. 3) MACL-IRFL incorporates and encodes the similarity relationships and the co-citation relationships into the representation learning of bug reports and source code files with weights, which helps to alleviate the insufficient bug-fix records issue and improve the accuracy of localization.
@@
To investigate whether the high-layer embedding propagation can help improve the performance of MACL-IRFL, we vary the number of layers $L$ of MACL-IRFL from 1 to 5 and compare their performance under different layers. Table 7 shows the overall performance of the five projects, and MACL-IRFL-$L$ denotes our model with $L$ layers embedding propagation. Fig. 9 illustrates the detailed performance of the five projects at different layer settings. We have the following observations.
@@ 8 Conclusion
[…]
Acknowledgements
This work was partially supported by National Key R&D Plan of China (Grant No.2024YFF0908003), Natural Science Foundation of China (No.62472326), CCF-Zhipu Large Model Innovation Fund (No.CCF-Zhipu202408), and the Major Program(JD) of Hubei Province (2023BAA018).